The cis-configuration of 2-butylene is higher in energy than the trans-configuration by 8.3E‑21J. You can synthesize trans-butylene with 100% yield. A customer wants to purchase your product but insists on 99.999% trans. What is the maximum temperature you can allow the product to be stored at to meet the customer’s specifications?
Answer #1

Answer:

We will use the Boltzmann Distribution law for the calculation;

p2/p1 = exp(-dE/KT). Here, we are assuming the degeneracies of two energy states are equal.

Here, dE = E2 - E1 = 8.3 x 10-21 J

Considering the p1 = 99.999 and p2 = 0.001

p2/p1 = 1.0001 x 10-5

and ln(p2/p1) = -dE/KT

T = -dE/K.ln(p2/p1) = -(8.3 x 10-21)/(1.380 x 10-23 x {-11.5128}) = 52.24 K

And maximum temperature should be 52.24 K

Please let me know, if you have any doubt by commenting below the answer.

Thanks